(1 point) Given the third order homogeneous constant coefficient equation y + 3y + 3y + y = 0 1) the auxiliary equation is
Answer #1

y + 39 + 3y + y = 0 m² +30m² +3m + 10 (m+ 13 = 0 m =-1,-1,-1 Three equal roots General Solution is y(x) = cente + Gaxen + C